OpenStudy (thomasfire55):

Solve for the roots in the following equation. (x^4 + 5x^2 - 36)(2x^2 + 9x - 5) = 0

zepdrix (zepdrix):

Where you stuck Tom Tom? Remember your Zero-Factor Property? \(\large\rm a\cdot b=0\qquad\to\qquad a=0\quad and/or\quad b=0\) When things multiply to give you zero, it means at least one of them is zero. Or in the case where you have a variable, it means both factors are zero. So this gives us our first step,\[\large\rm x^4+5x^2-36=0\qquad\qquad and\qquad\qquad 2x^2+9x-5=0\]

zepdrix (zepdrix):

Do you understand how to factor the second equation?